OpenStudy (anonymous):

a rectangular fish pond is 7ft longer than it it wide. a wooden 1ft wide is placed around the pound. the area covered by the pond and the walk is 58 sq ft greater than the area covered by the pond alone. what are the dimmensions of the pond?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

let length of pond be l and width = w 7ft longer than it it wide ----> l = w+7 got this ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

by placing 1 feet wide wooden, the total length and width increased by 1, right ? so, total length = l+1 total width = w+7+1 =w+8 got this ?

OpenStudy (anonymous):

yes i see

hartnn (hartnn):

the area covered by the pond and the walk is 58 sq ft greater than the area covered by the pond alone, area covered by the pond ----> lw = l (w+7) area covered by the pond and the walk ----> (l+1)(w+8) can you form an equation using this ?

hartnn (hartnn):

(l+1)(w+8) is 58 greater than l (w+7) so, what can be the equation ?

OpenStudy (anonymous):

1L-58=9w? what do i do

hartnn (hartnn):

how did you get that ? it'll be (l+1)(w+8) is 58 greater than l (w+7) so, if you add, 58 to l(w+7) they both will be equal, so, l(w+7) +58 = (l+1)(w+8) got this ?

hartnn (hartnn):

now just put there, l=w+7 and you get equation only in 'w'.

OpenStudy (anonymous):

50=1L+1w

hartnn (hartnn):

oh, sorry ,my typing mistake, it'll be (w+1)(w+8) is 58 greater than w(w+7) so, if you add, 58 to w(w+7) they both will be equal, so, w(w+7) +58 = (w+1)(w+8) now this is already in 'w' and the good thing is w^2 gets cancelled from both sides...

hartnn (hartnn):

w(w+7) +58 = (w+1)(w+8) w^2+7w+58 = w^2 +9w +8 7w+58 = 9w +8 isn't this easy to solve....

OpenStudy (anonymous):

25

hartnn (hartnn):

w= 25 is correct :) so. l=..?

OpenStudy (anonymous):

so whats the original cordinates

hartnn (hartnn):

l = w+7 =... ?

OpenStudy (anonymous):

32

hartnn (hartnn):

yes, so whats the original cordinates

OpenStudy (anonymous):

32

hartnn (hartnn):

means length = 32, width = 25

hartnn (hartnn):

these are the dimentions of the pond.

OpenStudy (anonymous):

57

hartnn (hartnn):

don't add, dimensions are length X width = 32 X 25

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome ^_^